Urmama182828:

2x^3+x^2+4 by x-3 Use polynomial long division show word

Manny300303199:

To divide the polynomial 2x^3 + x^2 + 4 by x - 3 using polynomial long division, follow these steps: Step 1: Write the division as a fraction using the divisor (x - 3) as the denominator: (2x^3 + x^2 + 4)/(x - 3) Step 2: Divide the leading term of the numerator (2x^3) by the leading term of the divisor (x): 2x^3 / x = 2x^2 Step 3: Multiply the divisor (x - 3) by the quotient from step 2: (x - 3) * 2x^2 = 2x^3 - 6x^2 Step 4: Subtract the result from step 3 from the original numerator: (2x^3 + x^2 + 4) - (2x^3 - 6x^2) = 7x^2 + 4 Step 5: Repeat steps 2-4 with the new numerator (7x^2 + 4): 7x^2 / x = 7x (x - 3) * 7x = 7x^2 - 21x (7x^2 + 4) - (7x^2 - 21x) = 25x + 4 Step 6: Repeat steps 2-4 with the new numerator (25x + 4): 25x / x = 25 (x - 3) * 25 = 25x - 75 (25x + 4) - (25x - 75) = 79 Step 7: The resulting polynomial 79 is the remainder. Therefore, when dividing 2x^3 + x^2 + 4 by x - 3 using polynomial long division, the quotient is 2x^2 + 7x + 25 and the remainder is 79.

dkononenko072:

First, set it up like a long division problem: markdown Copy code 2x^2 + 7x + 21 _________________________ x - 3 | 2x^3 + x^2 + 4 Now, divide 2 � 3 2x 3 by � x, which gives 2 � 2 2x 2 , so write 2 � 2 2x 2 above the line, then multiply � − 3 x−3 by 2 � 2 2x 2 and write it beneath the polynomial: markdown Copy code 2x^2 + 7x + 21 _________________________ x - 3 | 2x^3 + x^2 + 4 2x^3 - 6x^2 Now, subtract 2 � 3 − 2 � 3 2x 3 −2x 3 to get 0 0, and � 2 − ( − 6 � 2 ) x 2 −(−6x 2 ) to get 7 � 2 7x 2 , so bring down the next term: markdown Copy code 2x^2 + 7x + 21 _________________________ x - 3 | 2x^3 + x^2 + 4 2x^3 - 6x^2 _______________ 7x^2 Now, divide 7 � 2 7x 2 by � x, which gives 7 � 7x, so write 7 � 7x above the line, then multiply � − 3 x−3 by 7 � 7x and write it beneath the polynomial: markdown Copy code 2x^2 + 7x + 21 _________________________ x - 3 | 2x^3 + x^2 + 4 2x^3 - 6x^2 _______________ 7x^2 + 21x Now, subtract 7 � 2 − 7 � 2 7x 2 −7x 2 to get 0 0, and � 2 − 21 � x 2 −21x to get 21 � 21x, so bring down the next term: markdown Copy code 2x^2 + 7x + 21 _________________________ x - 3 | 2x^3 + x^2 + 4 2x^3 - 6x^2 _______________ 7x^2 + 21x 7x^2 - 21x Now, subtract 7 � 2 − 7 � 2 7x 2 −7x 2 to get 0 0, and 4 − ( − 21 � ) 4−(−21x)

curriful:

@dkononenko072 wrote:
First, set it up like a long division problem: markdown Copy code 2x^2 + 7x + 21 _________________________ x - 3 | 2x^3 + x^2 + 4 Now, divide 2 � 3 2x 3 by � x, which gives 2 � 2 2x 2 , so write 2 � 2 2x 2 above the line, then multiply � − 3 x−3 by 2 � 2 2x 2 and write it beneath the polynomial: markdown Copy code 2x^2 + 7x + 21 _________________________ x - 3 | 2x^3 + x^2 + 4 2x^3 - 6x^2 Now, subtract 2 � 3 − 2 � 3 2x 3 −2x 3 to get 0 0, and � 2 − ( − 6 � 2 ) x 2 −(−6x 2 ) to get 7 � 2 7x 2 , so bring down the next term: markdown Copy code 2x^2 + 7x + 21 _________________________ x - 3 | 2x^3 + x^2 + 4 2x^3 - 6x^2 _______________ 7x^2 Now, divide 7 � 2 7x 2 by � x, which gives 7 � 7x, so write 7 � 7x above the line, then multiply � − 3 x−3 by 7 � 7x and write it beneath the polynomial: markdown Copy code 2x^2 + 7x + 21 _________________________ x - 3 | 2x^3 + x^2 + 4 2x^3 - 6x^2 _______________ 7x^2 + 21x Now, subtract 7 � 2 − 7 � 2 7x 2 −7x 2 to get 0 0, and � 2 − 21 � x 2 −21x to get 21 � 21x, so bring down the next term: markdown Copy code 2x^2 + 7x + 21 _________________________ x - 3 | 2x^3 + x^2 + 4 2x^3 - 6x^2 _______________ 7x^2 + 21x 7x^2 - 21x Now, subtract 7 � 2 − 7 � 2 7x 2 −7x 2 to get 0 0, and 4 − ( − 21 � ) 4−(−21x)
mathway?

Manny300303199:

@curriful wrote:
@dkononenko072 wrote:
First, set it up like a long division problem: markdown Copy code 2x^2 + 7x + 21 _________________________ x - 3 | 2x^3 + x^2 + 4 Now, divide 2 � 3 2x 3 by � x, which gives 2 � 2 2x 2 , so write 2 � 2 2x 2 above the line, then multiply � − 3 x−3 by 2 � 2 2x 2 and write it beneath the polynomial: markdown Copy code 2x^2 + 7x + 21 _________________________ x - 3 | 2x^3 + x^2 + 4 2x^3 - 6x^2 Now, subtract 2 � 3 − 2 � 3 2x 3 −2x 3 to get 0 0, and � 2 − ( − 6 � 2 ) x 2 −(−6x 2 ) to get 7 � 2 7x 2 , so bring down the next term: markdown Copy code 2x^2 + 7x + 21 _________________________ x - 3 | 2x^3 + x^2 + 4 2x^3 - 6x^2 _______________ 7x^2 Now, divide 7 � 2 7x 2 by � x, which gives 7 � 7x, so write 7 � 7x above the line, then multiply � − 3 x−3 by 7 � 7x and write it beneath the polynomial: markdown Copy code 2x^2 + 7x + 21 _________________________ x - 3 | 2x^3 + x^2 + 4 2x^3 - 6x^2 _______________ 7x^2 + 21x Now, subtract 7 � 2 − 7 � 2 7x 2 −7x 2 to get 0 0, and � 2 − 21 � x 2 −21x to get 21 � 21x, so bring down the next term: markdown Copy code 2x^2 + 7x + 21 _________________________ x - 3 | 2x^3 + x^2 + 4 2x^3 - 6x^2 _______________ 7x^2 + 21x 7x^2 - 21x Now, subtract 7 � 2 − 7 � 2 7x 2 −7x 2 to get 0 0, and 4 − ( − 21 � ) 4−(−21x)
mathway?
seems so

Manny300303199:

@elsa213

shelovejayy:

\[\sqrt[htttt5667^{78}]{?}\]